OpenStudy (anonymous):

solve by mathematical induction : (6^n+2+7^2n+1)/43

OpenStudy (anonymous):

what is there to solve?

OpenStudy (anonymous):

\[\Large 43\left|(6^{n+2}+7^{2n+1})\right.\]

OpenStudy (anonymous):

provide me the solution to solve this problem dude

OpenStudy (anonymous):

Someone's in a rush.... LOL The solution, first, show that 43 divides that stuff when n = 1. That is... \[\Large 6^{1+2}+7^{2(1)+1}=6^3+7^3\]

OpenStudy (anonymous):

i want solution after substituting n=k dude

OpenStudy (anonymous):

Well, we assume that 43 does indeed divide \[\Large 6^{k+2}+7^{2k+1}\]

OpenStudy (anonymous):

Now all that's left to find is whether or not it divides the stuff when n = k+1

OpenStudy (anonymous):

thats the solution i needed

OpenStudy (anonymous):

\[\Large 6^{k+1+2}+7^{2(k+1)+1}=6^{k+3}+7^{2k+3}\]

OpenStudy (anonymous):

that's a stomper...

OpenStudy (anonymous):

i want the solution of proving that it is equal dude

OpenStudy (anonymous):

dude end the solution

OpenStudy (anonymous):

I'm still thinking :)

OpenStudy (anonymous):

k dude try to post after completing the solution

OpenStudy (anonymous):

besides, you should be throwing in your own ideas :P

OpenStudy (anonymous):

Let's try factoring a bit \[\Large 6\cdot 6^{k+2}+7^2\cdot 7^{2k+1}\] should be easy from here :)

OpenStudy (anonymous):

still blank? ^.^

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I uploaded the solution here, but first, I want to see your work :) http://www.megafileupload.com/en/file/442537/proof-43-divides-docx.html

terenzreignz (terenzreignz):

weird...